The Salesman Company currently sells its product for P15 and has variable costs of P7 per unit. Fixed costs are P50,000. Answer the following questions, consider each independently.

  • What is the breakeven point in units?
  • If fixed costs rise by 20%, what is the breakeven point in pesos?
  • If variable costs decline to P5, what is the breakeven point in units?
  • If the selling price per unit declines by 10%, what is the breakeven point in pesos?
1
Expert's answer
2021-08-11T09:14:11-0400

What is the breakeven points in units


"Breakeven point= \frac{fixed cost}{construction margin per unit}\\
=\frac{50000}{8}

= 6250"

If fixed costs rise by 20%. What is the breakeven point in units pesos?

New fixed cost = 50000

(+) 20% rise = 10000

60000

"Construction margin ratio= \frac {construction margin per unit} {selling price per unit}"

8/15*100= 53.33

Breakeven point in pesos= 60000/53.33*100= 112500

If variable costs decline to p5. What is the breakeven point in units?

Selling price per unit p 15

Variable cost per unit p 5

Construction margin per unit p 10

Breakeven point in units = 50000/10 = 5000